An electric bulb company manufacturing company claims that the bulbs produced by them have a mean life of 1250 hours. A random sample of 25 bulbs gives an average life of 1290 hours with standard deviation of of 110 hours. Use 5% level of significance and test the claim if bulb life follows is 1250 hours.


1
Expert's answer
2021-05-19T16:09:38-0400

1) H0: "\mu" =1250 (claim)

H1: "\mu" "\neq" 1250

2) Since "\alpha" =0.05 and the test is a two-tailed test, the critical values are -1.96 and +1.96

3)

"z=\frac{\overline{X}-\mu}{\sigma/\sqrt{n}}=\frac{1290-1250}{110/\sqrt{25}}=1.82"

4) Since -1.96<z<1.96, the decision is to not reject the null hypothesis.

5) There is enough evidence to support the claim.